OpenStudy (anonymous):

the surface area of a sphere is 100 square inches. use the formula s=4-πr^2 to find the length of the radius.

OpenStudy (mertsj):

\[100=4 \pi r^2\] \[\frac{25}{\pi}=r^2\] \[\frac{5}{\sqrt{\pi}}=r\]